# GenToC: Leveraging Partially-Labeled Data for Product Attribute-Value Identification

### v1

@@ 4.3. GenToC: Generative Attribute Extraction + Token Classification Value Extraction
Our proposed model, GenToC, operates as a two-step pipeline. The initial step employs a Generative Attribute Extraction (Gen-AE) model, which takes the product name as input and subsequently generates a concatenated list of predicted attributes. The model is trained to generate attributes in the order their values occur in the product name. Given the model’s generative ability, it might generate attributes outside the initial set. However, these are often relevant, so we forgo constrained generation (De Cao et al., 2021), avoiding any restrictions forcing the attributes to belong solely to the original set of attributes, .
@@
Markers: To deal with partially-labeled training data, we initially identify the words within the product name that correspond to values of any attribute. We refer to these identified words as marked words. For the example in Table 1, [boat, rockerz, 255, pro, raging, red] are treated as marked words during training, as they have been labeled with some attribute. A special learnable embedding, termed as “marker embedding”, is added to the encoder’s final hidden states of every token of the marked words before being passed to the decoder. The same learnable embedding is shared across all tokens everywhere. The model is then able to learn that the output attributes, as observed in the training data, are a result of considering only the limited set of words in the input that have been marked. Consequently, the marker embeddings act as signals that instruct the model to focus on and incorporate the information from these marked words into its attribute generation process, while preserving the broader context of the product name.
@@
When it comes to inference, the marker embedding is applied to all words within the product name, nudging the model to generalize and output attributes that are relevant to any word present in the product name. This enhances the model’s capability to recognize and associate attributes with words like bluetooth and neckband, which did not have valid attribute-value pairs annotated and hence were not marked at training time.
@@
In the next stage, a Token Classification Value Extraction (ToC-VE) model takes each of the attribute names from the first stage along with the original product name, separated by a special delimiter, ¡sep¿. It then labels each token with a binary value, for yes or no, to denote whether it corresponds to a value or not for the chosen attribute. This model is trained independently of the first-stage model. Since attribute names are used in the training process, it allows the model to learn from closely related attributes which share similarities in the attribute names. This becomes particularly crucial when dealing with a noisy attribute name ontology, which may include different attribute names conveying the same property, such as Model Number and Model No.
@@
Value Pruning: When training the ToC-VE model with a set of attribute-value pairs, it always contains a value for every attribute. However, during the inference phase, it depends on attributes produced by the Gen-AE model. As a consequence, even when the Gen-AE model generates incorrect attributes, the ToC-VE model tends to assign some value to them. To counteract this, we supplement the ToC-VE training with additional data to identify instances where no correct value is present for a given attribute. We accomplish this by taking an existing attribute-value pair and deleting the value from the product name. Consequently, we ensure that the chosen attribute no longer appears in the product name, training the model to label all tokens as ‘NO’ values for the attribute. For example, if we remove the brand, ‘boat’, the example, ‘brand ¡sep¿ rockerz 255 pro raging red bluetooth neckband’, will have no value for the attribute brand. We term the generation of this kind of synthetic training data as Value Pruning. For the final set of attribute-value pairs yielded by the model, we exclude those attributes that do not have any associated values.
Architecture | Partially labeled | Redundant attributes | Response time |
ToC-AVE | ✗ | ✗ | fastest |
Gen-AVE | ✗ | ✓ | slow |
GenToC | ✓ | ✓ | slow |
@@
Therefore, the architecture of GenToC ensures effective attribute-value pair extraction, preventing error build-up in the pipeline, even in the face of partially labeled data and vast attribute sets that contain redundancies. The complete algorithm for training and inference are described in Algorithm 1 and Algorithm 2, respectively.
@@ 5.1. Dataset and Metrics
To curate data for the attribute-value extraction task, we make use of the product specifications that are provided by the sellers for product listings on IndiaMART111https://www.indiamart.com. IndiaMART is India’s largest online B2B marketplace, connecting 182M buyers with 7.7M suppliers and featuring over 102M different products and services. We use a set of 2M examples for training models in this work. It comprises 715K unique attribute-value pairs covering 24.7K attributes. The category-wise distribution of products is presented in Figure 2. Table 3 lists the top-five attributes and provides a set of example values for each. Despite the large scale of attributes, only 40.7% of the words of a product name are tagged on average with an attribute. This motivates us to build models designed for partially labeled data, where the present attribute-value pairs are reliable and of high quality but might lack the complete set of attribute-values.
@@
We evaluate the systems on 39,671 samples (39K) based on the ground truth values available. We compute the precision, recall and F1-score by comparing the set of attribute-value pairs generated for each input with the ground truth set of attribute-value pairs. We then report their averages taken across all examples. However, we find that automatic evaluation is challenging and often unreliable due to the lack of normalization in the attribute names, as the ground truth set can use different attributes to express the same characteristic. So we randomly sample 2,000 examples (2K) and get every output checked using 3 data annotators (DAs). These annotators are skilled professionals who perform various annotation tasks within our organization. We consider the majority class assigned by the DAs to determine if a given attribute-value pair is correct/incorrect. We find that the inter-annotator agreement, computed using Fleiss’ kappa (Fleiss, 1971), is . Response time is measured by taking the average on the same 2K queries using NVIDIA GeForce RTX 3090 GPU (with batch size of one).
@@ 5.2. Implementation and Systems Compared
To compare GenToC with the two classes of state-of-art models, we use ToC-AVE as a representative of the available NER-style models (Zheng et al., 2018; Rezk et al., 2019; Shinzato et al., 2022) and Gen-AVE as a representative of the published generative models (Roy et al., 2022; Shinzato et al., 2023). We use our own implementations due to lack of availability of standard open-source implementations for the above works. We use the DeBERTa-V3-Small222https://hf.co/microsoft/deberta-v3-small (He et al., 2021) for all the classification tasks and BART-Base333https://hf.co/facebook/bart-base (Lewis et al., 2019) for all generation tasks. Both architectures consist of 6 encoder layers, and BART has 6 additional decoder layers. Due to the need for industry-scale systems to be used by millions of users, we don’t experiment with LLMs, which are highly resource intensive.
@@
6. Experiments and Results
In this section, we answer the following questions:
What is the significance of Markers and Value Pruning in the GenToC system?
Which system is better when considering the trade-off between precision and recall?
How do the models perform when dealing with long product names?
Architecture | Data Source | Automatic Evaluation (39K) | Manual Evaluation (2K) | ||||
@@ 6. Experiments and Results
Product Name | ToC-AVE | Gen-AVE | GenToC |
@@ 6.1. Comparison with Other Models
In Table 4, we present a comparison of three systems – ToC-AVE, Gen-AVE and GenToC. GenToC model outperforms its counterparts in recall, achieving 71.8% in automatic and 80.1% in manual evaluations, which are the highest among the three models. These scores are notably higher (by 11.7% and 27.3% in automatic and manual evaluations) than those of the next best-performing model, ToC-AVE. Such a significant gap in recall proves that GenToC effectively leverages the partially labeled nature of the training data, resulting in more extractions. Indeed, the total number of correct attribute-value extractions rose from 2,636 to 4,121 (a 56.3% increase) on the 2K test set. In terms of precision, both the ToC-AVE and Gen-AVE models exhibit higher precision than the GenToC model in both automatic and manual evaluations, with approximately 9-10% and 4-5% higher precision, respectively.
Overall, GenToC maintains its superiority in F1-score, recording the highest at 71.3% in automatic evaluations and 83.0% in manual evaluations, indicating a more balanced and consistent performance. However, a drawback of GenToC is its slower response time compared to the ToC-AVE model. We address a potential solution to this performance-speed trade-off in Section 6.3.
In Table 6, we present example cases showcasing GenToC’s superior performance in comparison to both ToC-AVE and Gen-AVE models, in terms of the number of attribute-value pairs extracted. For instance, in the first example, the GenToC model surpasses the other models by identifying four additional attributes – Occasion, Sleeves Type, Pattern and Gender.
@@ 6.2. Ablation Study
We also experiment with adding Markers to ToC-AVE and Gen-AVE. During training, we incorporate marker embeddings to those words which have been covered by the available attributes, and at inference, we incorporate the marker embeddings to all words. Consistent with earlier practices, these embeddings are added into the final hidden states of the encoder. In the 2K test set, this modification results in considerable reductions in precision scores, plummeting to 66.5% from 90.8% in ToC-AVE and to 65.2% from 89.6% in Gen-AVE. A similar pattern can be observed in case of the 39K test set as well. This outcome underscores the importance of two-stage model for using markers effectively. The bifurcated structure of the two-stage model ensures that markers are exclusively utilized in the first stage. This design circumvents the problem in single-stage models, where markers can force the assignment of outputs to every word, often resulting in suboptimal performance.
@@ 6.3. Bootstrapping Training Data
From Table 4, we find that ToC-AVE has the fastest response time. Gen-AVE is the slowest, as it has to generate a long string containing all attributes and values. The high performance of GenToC comes at the cost of speed, as its response time is 10x greater than ToC-AVE’s. With the motivation to build a deployable system that can handle real-world traffic, we experiment with cleaning the training data by regenerating it using the trained GenToC model. It increases the average number of words tagged in any attribute-value pair from 0.41 to 0.65 while increasing the total number of attribute-value pairs from 3M to 4.7M on the full training data. Furthermore, it also brings down the number of attributes from 24.7K to 7.2K, potentially minimizing the presence of redundant and un-canonicalized attributes. In Table 5, we find that the ToC-AVE model trained with this GenToC-bootstrapped training data significantly improves the performance. In fact, its performance is comparable to the GenToC model in both automatic and manual evaluations. This makes it a strong alternative to the GenToC model for production environments, where rapid response times are crucial.
@@ 6.4. Precision-Recall Trade-off
To understand the trade-off between precision and recall across different systems, we evaluate the systems using a precision-recall curve. We employ a common re-scoring model to compute the confidence level for all system extractions in this process. The model used for re-scoring is an independent Seq2Seq model, trained by using the ‘product name’ as the input and the ‘attribute: value’ string as the output, utilizing the same training data. The model computes the confidence level associated with any given attribute-value pair by accumulating the log probabilities for every token present within the output string. We found that the model provides us with better-calibrated scores for an attribute-value pair generated by any of the systems. By applying uniformly spaced thresholds between 0 and 1 for the confidence values, we created the Precision-Recall curve shown in Figure 3. Upon inspection, it was clear that GenToC’s performance significantly eclipsed that of other models. Across most of the curve, GenToC achieves a higher precision for a given recall, demonstrating its power. This substantiates the superior quality and efficacy of the GenToC system.
6.5. Performance on Long Product Names
To assess how well the models handle intricate scenarios, we evaluate their performance on unusually long product names. Product names consist of five words on average, with a standard deviation of two. Thus, we create a test set comprising 500 examples, each containing at least seven words, for the evaluation process. In Table 7, we tabulate the scores obtained from manual evaluation, together with the tagged ratio — that is, the proportion of words in the product name that have been marked with an attribute — for various models.
While Gen-AVE exhibits the highest precision at 95.3%, and GenToC without VP shows a notable recall of 90.3%, GenToC, stands out with the highest F1-score of 90.7%. This underscores GenToC’s superior balance in the precision-recall trade-off, even in cases where the product names are long. It’s noteworthy that when Markers are incorporated into the ToC-AVE and Gen-AVE models, over 99% of the words in product names are linked to an attribute. However, this high tagging ratio might not be ideal, particularly for lengthy queries, as not every word necessarily possesses a relevant attribute. This is also reflected in the significant decrease in precision values for these models when Markers are used — a drop of over 22% compared to their counterparts without Markers. Single-stage models, which have to perform value labeling alongside attribute extraction, face this issue when Markers are employed, as they attempt to arbitrarily assign an attribute to every word. However, since we apply markers only to the first stage of the GenToC model, it does not face this problem as there is no direct one-to-one mapping between the predicted attributes and input words. Finally, even the ToC-AVE model trained with GenToC bootstrapping attains an F1-score that’s comparable with GenToC, while having a faster response time. This shows that the full potential of ToC-AVE model is realized only with high-quality training data.
Architecture | Precision | Recall | F1-score |
ToC-AVE | 94.8 | 65.2 | 77.3 | 0.464 | ||
with Marker | 72.4 | 90.3 | 80.4 | 0.999 | ||
Gen-AVE | 95.3 | 58.2 | 72.2 | 0.403 | ||
with Marker | 71.7 | 71.6 | 71.6 | 0.993 | ||
GenToC | 90.7 | 90.7 | 90.7 | 0.662 | ||
without Marker | 94.5 | 58.7 | 72.4 | 0.405 | ||
without VP | 87.1 | 93.0 | 89.9 | 0.697 | ||
without Marker & VP | 94.2 | 59.0 | 72.6 | 0.406 | ||
90.5 | 88.4 | 89.4 | 0.689 |

### v2

@@ 5. Methods
In this section, we describe the two-stage GenToC model and the marker-based training that enables it to learn from partially-labeled data. We show the working of GenToC in Figure 2 (c). The initial step employs a Generative Attribute Extraction (Gen-AE) model, which takes the product title or user search query as input and subsequently generates a concatenated list of predicted attributes. The model is trained to generate attributes in the order their values occur in the input. Given the model’s generative ability, it might generate attributes outside the training set. However, these are very infrequent (occur less than 0.1% of the times), so we forgo constrained generation (De Cao et al., 2021), avoiding any restrictions that would force the attributes to belong solely to the original set of attributes.
Markers:
@@
To deal with partially-labeled training data, we initially identify the words within the input query that correspond to values of any attribute. We refer to these identified words as marked words. For the example, “Boat Rockerz 255 Pro Raging Red Bluetooth Neckband”, in Table 1, [boat, rockerz, 255, pro, raging, red] are treated as marked words during training, as they have been labeled with some attribute. A special learnable embedding, termed as “marker embedding”, is added to the encoder’s final hidden states of every token of the marked words before being passed to the decoder. The same learnable embedding is shared across all marked tokens everywhere. The model is then able to learn that the output attributes, as observed in the training data, are a result of considering only the limited set of words in the input that have been marked. In other words, the marker embeddings act as signals that instruct the model to focus on and incorporate the information from these marked words into its attribute generation process, while preserving the broader context of the input query.
@@
At inference time, the marker embedding is applied to all words within the input query, nudging the model to generalize and output attributes that are relevant to any word present in the query. This enhances the model’s capability to recognize and associate attributes with words like bluetooth and neckband, which did not have valid attribute-value pairs annotated and hence were not marked at training time.
@@
In the next stage, a Token Classification Value Extraction (ToC-VE) model takes each of the attribute names from the first stage along with the original query, separated by a special delimiter, ¡sep¿. It then labels each token with a binary value, for yes or no, to denote whether it corresponds to a value or not for the chosen attribute. This model is trained independently of the first-stage model. Since attribute names are used in the training process, it allows the model to learn from closely related attributes which share similarities in their names. This becomes particularly crucial when dealing with a noisy attribute name ontology, which may include different attribute names conveying the same property, such as Model Number and Model No.
Value Pruning:
@@
When training the ToC-VE model with a set of attribute-value pairs, it always contains a value for every attribute. However, during the inference phase, the attributes may be erroneously generated by the Gen-AE model. As a consequence, ToC-VE model would tend to assign some value to even the incorrect attributes. To counteract this, we supplement the ToC-VE training with additional data to identify instances where no correct value is present for a given attribute. We accomplish this by taking an existing attribute-value pair and deleting the value from the original query. Consequently, we ensure that the chosen attribute no longer appears in the query, training the model to label all tokens as ‘NO’ values for the attribute. For example, if we remove the token, ‘boat’, the example, “brand ¡sep¿ rockerz 255 pro raging red bluetooth neckband”, will have no value for the attribute brand. We term the generation of this kind of synthetic training data as Value Pruning. For the final set of attribute-value pairs yielded by the model, we exclude those attributes that do not have any associated values.
@@
Therefore, the architecture of GenToC ensures effective attribute-value pair extraction, preventing error build-up in the pipeline, even in the face of partially labeled data and vast attribute sets that contain redundancies. The complete algorithm for training and inference are described in Algorithm 1 and Algorithm 2 (in Appendix), respectively.
@@ 6.1. Dataset and Metrics
To curate data for the attribute-value extraction task, we make use of the product specifications that are provided by the sellers for product listings on IndiaMART. IndiaMART has 194M buyers, 7.9M sellers, and features over 108M different products and services. For the experiments in this paper, we limit our input exclusively to product titles and use a set of 2M examples for training models. It comprises 715K unique attribute-value pairs covering 24.7K attributes. The category-wise distribution of products is presented in Figure 3. Table 3 lists the top-five attributes and provides a set of example values for each. Despite the large scale of attributes, only 40.7% of the words of a product title are tagged on average with an attribute. This motivates us to build models designed for partially labeled data, where the present attribute-value pairs are reliable and of high quality but might lack the complete set of attribute-values.
@@
We evaluate the systems on 39,671 samples (39K) based on the ground truth values available. We compute the precision, recall and F1-score by comparing the set of attribute-value pairs generated for each input with the ground truth set of attribute-value pairs. We then report their averages taken across all examples. However, we find that automatic evaluation is challenging and often unreliable due to the lack of normalization in the attribute names, as the ground truth set can use different attributes to express the same characteristic. So, we randomly sample 2,000 examples (2K) and get every output checked using 3 data annotators (DAs). These annotators are skilled professionals who perform various annotation tasks within our organization. We consider the majority class assigned by the DAs to determine if a given attribute-value pair is correct/incorrect. We find that the inter-annotator agreement, computed using Fleiss’ kappa (Fleiss, 1971), is , indicating substantial agreement among the annotators (Landis and Koch, 1977). The response time is measured by calculating the average time taken for queries from the 2K test set (with a batch size of one) using an NVIDIA GeForce RTX 3090 GPU.
@@ 6.2. Implementation and Systems Compared
To compare GenToC with the two classes of state-of-art models, we use NER-AVE as a representative of the available NER-style models (Zheng et al., 2018; Rezk et al., 2019; Shinzato et al., 2022) and Seq2Seq-AVE as a representative of the published generative models (Roy et al., 2022; Shinzato et al., 2023). We use our own implementations due to the lack of availability of standard open-source implementations for the above works. For a fair comparison, we use the DeBERTa-V3-Small222https://hf.co/microsoft/deberta-v3-small (He et al., 2021) for all the classification tasks and BART-Base333https://hf.co/facebook/bart-base (Lewis et al., 2019) for all generation tasks. Both architectures consist of 6 encoder layers, and BART has 6 additional decoder layers. Due to the need for industry-scale systems to be used by millions of users, we don’t experiment with LLMs, which are highly resource intensive.
7. Experiments and Results
In this section, we answer the following questions:
What is the incremental contribution of Markers and Value Pruning in the GenToC system?
What are the trade-offs between precision and recall in various models at different confidence thresholds?
How is the system utilized and how does it perform in a production environment?
@@ 7. Experiments and Results
Data Source | Architecture | Automatic Evaluation (39K) | Manual Evaluation (2K) | ||||
@@
Product Title | NER-AVE | Seq2Seq-AVE | GenToC |
@@ 7.1. Comparison with Other Models
In Table 4, we present a comparison of three systems – NER-AVE, Seq2Seq-AVE and GenToC. GenToC model outperforms its counterparts in recall, achieving 71.8% in automatic and 80.1% in manual evaluations, which are the highest among the three models. These scores are notably higher (by 11.7% and 27.3% in automatic and manual evaluations) than those of the next best-performing model, NER-AVE. Such a significant gap in recall demonstrates that GenToC effectively leverages the partially labeled nature of the training data, resulting in more extractions. Indeed, the total number of correct attribute-value extractions rose from 2,636 to 4,121 (a 56.3% increase) on the 2K test set. This increase in recall is mainly due to Markers, which is elaborated upon in Section 7.2. In terms of precision, both the NER-AVE and Seq2Seq-AVE models exhibit higher precision than the GenToC model in both automatic and manual evaluations, with approximately 9-10% and 4-5% higher precision, respectively. Overall, GenToC maintains its superiority in F1-score, recording the highest at 71.3% in automatic evaluations and 83.0% in manual evaluations, indicating a more balanced and consistent performance. However, a drawback of GenToC is its slower response time compared to the NER-AVE model. We address a potential solution to this performance-speed trade-off in Section 7.3.
In Table 6, we present example cases showcasing GenToC’s superior performance in comparison to both NER-AVE and Seq2Seq-AVE models, in terms of the number of attribute-value pairs extracted.
NER-AVE and Seq2Seq-AVE models are unable to leverage the partially-labeled nature of the training data, as indicated previously in Table 2, despite possibly encountering the remaining unlabeled attribute-value pairs in other examples. In particular, each training example is tagged with only 1.53 attribute-value pairs on average, so it’s not unexpected that these models maintain a similar tagging frequency during inference. GenToC addresses this challenge by incorporating Markers, which allow for the identification of more number of attribute names than what is typically found in the training data.
For instance, in the product title “Casual Juliet Sleeve Solid Women Maroon Top” from the first example in Table 6, the GenToC model surpasses the other models by identifying four additional attributes – Occasion, Sleeves Type, Pattern and Gender.
@@ 7.2. Ablation Study
We also experiment with adding Markers to NER-AVE and Seq2Seq-AVE. During training, we incorporate marker embeddings to those words which have been covered by the available attributes, and at inference, we incorporate the marker embeddings to all words. Consistent with earlier practices, these embeddings are added into the final hidden states of the encoder. In the 2K test set, this modification results in considerable reductions in precision scores, plummeting to 66.5% from 90.8% in NER-AVE and to 65.2% from 89.6% in Seq2Seq-AVE. A similar pattern can be observed in case of the 39K test set as well. This outcome underscores the importance of two-stage model for using markers effectively. The bifurcated structure of the two-stage model ensures that markers are exclusively utilized in the first stage. This design circumvents the problem in single-stage models, where markers can force the assignment of attributes to every word, often resulting in suboptimal performance. This also significantly increases the response time for the Seq2Seq-AVE model (see Table 4), as it has to generate a longer output due to more extractions. On the other hand, Value Pruning is useful when there is an candidate set of attribute names that might include some incorrect ones, such as those generated by the first stage of the GenToC model. It aims to reduce irrelevant attribute names in the final result by predicting empty values for them. Conversely, Seq2Seq-AVE and NER-AVE produce both attribute-value pairs in a unified step, making Value Pruning inapplicable for them.
@@ 7.3. Bootstrapping Training Data
From Table 4, we find that NER-AVE has the fastest response time. Seq2Seq-AVE is the slowest, as it has to generate a long string containing all attributes and values. The high performance of GenToC comes at the cost of speed, as its response time is 10x greater than NER-AVE’s. With the motivation to build a deployable system that can handle real-world traffic, we experiment with cleaning the training data by regenerating it using the trained GenToC model. It increases the average number of words tagged in any attribute-value pair from 0.41 to 0.65 while increasing the total number of attribute-value pairs from 3M to 4.7M on the full training data. In Table 5, we find that the NER-AVE model trained with this GenToC-bootstrapped training data significantly improves the performance over the NER-AVE model trained with original partially-labeled data. In fact, its performance when trained with GenToC-bootstrapped data is comparable to the GenToC model in both automatic and manual evaluations. This makes it a strong alternative to the GenToC model for production environments, where rapid response times are crucial. Additionally, we experiment with regenerating the training data using the originally trained NER-AVE, the second-best model in terms of F1-score. However, we find that training with this data results in marginally decreased performance, proving less effective than training with GenToC-bootstrapped data. This is because, unlike GenToC, the NER-AVE model is not equipped to learn from incomplete data. Consequently, it does not help in reducing the partially labeled nature of the training data upon regeneration, rendering the bootstrapping unproductive. Furthermore, in Appendices B and C, we also assess the performance of these models on long product titles and long-tail attributes. We find that GenToC outperforms Seq2Seq-AVE and NER-AVE systems by more than 13-18% in F1-score for long product titles and 7-19% in precision for rare attributes while helping boostrapped NER-AVE achieve similar levels.
@@ 7.4. Precision-Recall Trade-off
To understand the trade-off between precision and recall across different systems, we evaluate the systems using a precision-recall curve. We employ a common re-scoring model to compute the confidence level for all system extractions in this process. The model used for re-scoring is an independent Seq2Seq model, trained by using the product title as the input and the ‘attribute: value’ string as the output, utilizing the same training data. The model computes the confidence level associated with any given attribute-value pair by accumulating the log probabilities for every token present within the output string. We found that the model provides us with better-calibrated scores for an attribute-value pair generated by any of the systems. Using the results from the manually evaluated (2K) set, we applied uniformly spaced thresholds between 0 and 1 for the confidence values to create the Precision-Recall curve shown in Figure 4. Upon inspection, it was clear that GenToC’s performance significantly eclipsed that of baselines. Across most of the curve, GenToC achieves a higher precision for a given recall, demonstrating its power. Further, the AUC in case of NER-AVE and Seq2Seq-AVE are 0.51 and 0.48 respectively, while that of GenToC is 0.76. This substantiates the superior quality and efficacy of the GenToC system.
Additionally, the curve for NER-AVE trained with GenToC bootstrapped data closely mirrors that of GenToC, with an AUC of 0.77. This further demonstrates the effectiveness of training with data bootstrapping from GenToC.
7.5. Deployment Status and Impact
Our system has been deployed on India’s largest B2B e-commerce platform, IndiaMART, with 194M buyers and 7.9M sellers. It has been successfully integrated into the core product search functionality and has already served over 200Mrequests since deployment. Specifically, it extracts attribute-value pairs from product titles in listings, and enables dynamic feature highlighting of product features according to user search queries. As an illustration of dynamic feature highlighting, when a user searches for “10 tier shoe rack”, the system detects the attribute-value pair Number of shelves: 10. This information is then used to highlight the corresponding feature Available number of shelves: 10 available for a listing with the product title “Steel Shoe Rack”, in the search results. This leads to enriched user search experience.
For the deployment version, we train GenToC using a dataset that is 3x larger, with a similar partially-labeled nature. Further, we regenerate the training dataset using the trained GenToC model and subsequently train a faster NER-AVE model with this bootstrapped data, as detailed in Section 7.3. We use this GenToC-bootstrapped NER-AVE model for deployment.
System | Precision | Recall | F1-score |
Rule-based (prior deployment) | 91.9 | 70.9 | 80.0 |
Proposed method | 89.5 | 91.1 | 90.3 |
We find significant improvements over the prior deployed system (which is based on rule-based non-neural system) for this task. Rule-based systems that use regular-expression based techniques have their own limitations, such as struggling with negations (e.g., non-stretchable jeans), handling common spelling variations (litre, liter, l, or ltr), making broad generalizations (both red and raging red refer to colors), and properly tagging attributes contextually (Galaxy could refer to either a chocolate brand or a Samsung mobile model).
In an offline evaluation, based on a manual audit of search queries, we observe the results tabulated in Table 7. It can be observed that the previously deployed rule-based system falls short of our proposed approach by more than 20% in terms of recall, while being only marginally better by less than 2.5% in precision. This results in an overall difference of approximately 10% in the F1-score between the two methods. In online evaluation, we find that our system leads to a 9% increase in queries with identified attributes that lead to dynamic feature highlighting.